1 - Yields on municipal bonds with tax benefit are typically ________ yields on corporate bonds of similar risk and time to maturity.

Group of answer choices

lower than

twice as high as

identical to

slightly higher than

2-

When a coupon bond matures, the issue pays the principal but not the coupon.

Group of answer choices

True

False

3-

Price a 2-yr 5% semiannual coupon bond with a par value of $1000 and the yield to maturity (discount rate) is 4.84% per annum.

Round the answer to the nearest 2 decimal points. If your answer is $12.345, enter "12.35" in the answer box.
Answer #1

1)

lower than

Since municipal bonds provide tax benefits, the yield to maturity which is basically the return to investors will be lower than corporate bonds.

2)

FALSE

A maturity, a bond will typically pay principal and the final coupon.

3)

Rate = 4.84% / 2 = 2.42%

Coupon = (0.05 * 1000) / 2 =

Number of periods = 2 * 2 = 4

Price = Coupon * [1 - 1 / (1 + r)n] / r + FV / (1 + r)n

Price = 25 * [1 - 1 / (1 + 0.0242)4] / 0.0242 + 1000 / (1 + 0.0242)4

Price = 25 * 3.769235 + 908.784506

Price = $1,003.02